# Beamforming and Waveform Optimization for RF Wireless Power Transfer with Beyond Diagonal Reconfigurable Intelligent Surfaces

### v1

@@ IV-A Convergence Analysis
Fig. 2 illustrates the convergence performance of the IT-BDRIS, IT-WF, iterative passive beamforming, and SDR-BDRIS. As seen in Fig. 2.a and Fig. 2.b, both beamforming and waveform optimization algorithms iteratively converge toward a local optimum solution. Moreover, the number of required iterations increases with $M$ and $N$ as discussed in Section III-D. Moreover, Fig. 2.c and Fig. 2.d highlight that the proposed methods for alternating optimization of waveform and passive beamforming lead to convergence. Although the number of iterations in SDR-BDRIS is lower than in IT-BDRIS, each iteration of SDR-BDRIS requires solving multiple SDP problems with high complexity. At the same time, IT-BDRIS relies only on closed-form computations. Observe that the number of iterations for different setups is the same in SDR-BDRIS, but the time complexity of solving (25) drastically increases with $M$ since the hermitian matrix sub-space scales with $M^{4}$.
@@ IV-B Algorithms Performance Comparison
Herein, we compare the performance of the proposed algorithms and consider also an optimal benchmark derived for $N=1$. Note that for a single-carrier system ($N=1$), problem (5) becomes equivalent to maximizing the RF power, i.e., the second-order term in $i_{dc}$, in terms of the solution. Thus, the framework introduced in [36] serves as a benchmark when $N=1$ and is labeled as optimal in Fig. 3.a.
Fig. 3 compares the performance of the proposed algorithms for $N=1$ and $N=4$ as a function of $M$. Note that SDR-BDRIS achieves optimal performance, while IT-BDRIS leads to relatively lower $i_{dc}$ values, but is still close to the optimal solution and the gap increases with $M$. Moreover, for $N=4$, the same trend is observed, where SDR-BDRIS outperforms IT-BDRIS. However, as stated earlier, the complexity of SDR-BDRIS is considerably higher than that of IT-BDRIS, especially for large $M$; thus, a trade-off exists between performance and complexity. Note that there is no optimal benchmark for the multi-carrier case. Fig. 4 compares the performance of the proposed algorithms as a function of $N$, where SDR-BDRIS outperforms IT-BDRIS also in this case.
Figure 3: Average $i_{dc}$ at the ER as a function of $M$ using different algorithms for (a) $N=1$ (left) and (b) $N=4$ (right) with NLoS channel ($\alpha=0.1$), $P_{T}=50$ dBm, and a fully-connected BD-RIS.
Figure 4: Average $i_{dc}$ at the ER as a function of $N$ using different algorithms with $\alpha=0.1$, $M=16$ and $P_{T}=50$ dBm, and a fully-connected BD-RIS .
@@ IV-C Waveform Behaviour & Channel Analysis
One of the key considerations of WPT with non-linear EH is the operating regime of the rectifier. Specifically, if the input RF power to the rectifier is relatively low, the RF-to-DC conversion efficiency drops significantly. On the other hand, if the RF input power is high such that the rectifier’s diode enters the breakdown region, the mathematical model for the small-signal regime is no longer valid. Thus, it is important to select $P_{T}$ properly such that the received RF power is in the proper regime [8]. For this, we proceed with some discussions on the system’s behavior for different $P_{T}$ values.
@@
Figure 5: The normalized gain of the channels and input signal at different sub-carriers for (a) $P_{T}=50$ dBm (top) and (b) $P_{T}=30$ dBm (bottom) for $\alpha=0.1$ (left), $\alpha=1$ (middle), and $\alpha=10$ (right). We assume a fully-connected BD-RIS with $M=32$, $N=8$, and a pure NLoS ($\kappa=0$) random channel realization. Note that the bars are represented in a stacked fashion, meaning that bars with the same length represent the same normalized gain.
@@
Figure 6: The received signal in time domain for (a) $P_{T}=50$ dBm (top) and (b) $P_{T}=30$ dBm (bottom) for $\alpha=1$ (left) and $\alpha=10$ (right) with fully-connected BD-RIS, $M=32$, and $N=8$ at a random channel realization.
@@ IV-D D-RIS versus BD-RIS
Herein, we compare the performance of D-RIS and fully connected BD-RIS. The BD-RIS results are obtained using the SDR-BDRIS approach, while for D-RIS, this approach can be straightforwardly modified as presented in [34].
Note that for far-field LoS, the channel is deterministic and is computed using the relative phase difference of elements. Thus, we assume that the variation among the sub-carriers is negligible such that for a given $\boldsymbol{\Theta}$, $h_{n}=h$. Hereby, (II-B) can be rewritten as $\alpha_{1}\left\lVert h\right\rVert^{2}+\alpha_{2}\left\lVert h\right\rVert^{4}$, where $\alpha_{1}$ and $\alpha_{2}$ are positive coefficients depending on the waveform and rectenna circuit. To maximize this, it is sufficient to maximize the cascade channel gain $\left\lVert h\right\rVert$, which can be achieved by constructing a diagonal $\boldsymbol{\Theta}$ such that $[\boldsymbol{\Theta}]_{i,i}=e^{-j([\tilde{\mathbf{h}}_{R,n}]_{i}+[\tilde{\mathbf{h}}_{I,n}]_{i})},\forall i$. Such a scattering matrix can be designed using D-RIS; thus, D-RIS can achieve the optimal performance and BD-RIS does not introduce any additional gain. Moreover, for $N=1$, (5) becomes equivalent to maximizing the RF power, which allows for extending the findings of [16] to single-carrier WPT. Thus, for a channel with NLoS components, we expect that BD-RIS outperforms D-RIS even with a continuous wave.
@@
V Conclusion
We considered a BD-RIS-aided SISO WPT system with EH non-linearity. Moreover, we formulated a joint beamforming and waveform optimization problem aiming to maximize the harvested power at the ER. To address this problem, we decoupled the waveform optimization and beamforming problems and leveraged alternating optimization. We proposed an efficient iterative method for waveform optimization and three beamforming approaches namely SDR-BDRIS, SDP-BDRIS, and IT-BDRIS. Furthermore, IT-BDRIS has the lowest computational complexity but in cost of performance losses. Meanwhile, It was seen that SDR-BDRIS achieves the same performance as SDP-BDRIS with less computational complexity. The simulation results proved that the proposed algorithms converge to a local optimal solution, while the complexity scales with the number of RIS elements and the number of sub-carriers. Moreover, SDR-BDRIS achieves the optimal performance for a single-carrier system and outperforms IT-BDRIS in other scenarios. Interestingly, we proved that BD-RIS achieves the same performance as D-RIS under LoS conditions (in the absence of mutual coupling), while it outperforms D-RIS in NLoS cases. Furthermore, our findings demonstrated that the frequency selectivity of the channel impacts the way that BD-RIS shapes the cascade channel. Finally, we showed that the waveform power allocation to different sub-carriers depends on the transmit power and the rectifier’s operating regime.
Appendix A Proof of Theorem 1
Let us proceed by writing the Lagrangian of (6) as
By setting (41) to zero, we attain (11). Moreover, the optimal solution must utilize the whole transmit power budget, i.e., $\frac{1}{2}\sum_{n=1}^{N}{\bar{s}_{n}}^{2}=P_{T}$. Thus, by substituting (11) into the latter equality, the dual variable $\lambda^{(l)}$ is given by (12). Finally, we introduce an additional parameter $\rho_{s}\leq 1$ to control the convergence of the solution such that the subgradient update is given by (10). ∎
Appendix B Proof of Proposition 1
@@
=$\stackrel{{\scriptstyle\textnormal{(a)}}}{{\mathstrut{=}}}\mathrm{Tr}({\mathbf{h}_{I,n}}\mathbf{h}_{R,n}^{T}\boldsymbol{\Theta})=\mathrm{Tr}(\mathbf{H}_{n}\boldsymbol{\Theta})$
=$\stackrel{{\scriptstyle\textnormal{(b)}}}{{\mathstrut{=}}}{\mathrm{Vec}(\mathbf{H}_{n})}^{T}\mathrm{Vec}(\boldsymbol{\Theta})\stackrel{{\scriptstyle\textnormal{(c)}}}{{\mathstrut{=}}}{\mathrm{Vec}(\mathbf{H}_{n})}^{T}\mathbf{P}\boldsymbol{\theta}=\mathbf{a}_{n}^{T}\boldsymbol{\theta}.$ (42)
Assume that $\mathbf{D}$, $\mathbf{F}$, and $\mathbf{H}$ are arbitrary matrices. Hereby, (B) and (42) come from $\mathrm{Tr}(\mathbf{DFH})=\mathrm{Tr}(\mathbf{HDF})$ and $\mathrm{Tr}(\mathbf{D}^{T}\mathbf{F})=\mathrm{Vec}(\mathbf{D})^{T}\mathrm{Vec}(\mathbf{F})$, respectively. Moreover, by defining $\boldsymbol{\theta}\in\mathbb{C}^{\bar{M}\times 1}$ as the vector containing the lower/upper-triangle elements in $\mathbf{\Theta}$, one can design a permutation matrix $\mathbf{P}\in\{0,1\}^{M^{2}\times\bar{M}}$ such that $\mathbf{P}\boldsymbol{\omega}=\mathrm{Vec}(\mathbf{\Omega})$ holds, which leads to (42). ∎
Appendix C Proof of Proposition 2
The Neumann approximation can be used to approximate the inverse of a matrix using another close matrix. Imagine $\mathbf{A}$ is a close matrix to an invertible $\mathbf{X}$, such that [46]
@@ Appendix C Proof of Proposition 2
Then, the inverse of $\mathbf{A}$ can be written using the Neumann series as [47]
@@
Now, let us consider $\mathbf{\Omega}$ as a small increment to the impedance matrix $\mathbf{Z}$, such that $\left\lVert[\mathbf{\Omega}]_{i,m}\right\rVert\leq\delta,\forall i,m$, leading to $\mathbf{Z}+\mathbf{\Omega}$ being close to $\mathbf{Z}$. Hereby, we can write the first-order Neumann series as in (27). Note that the first-order approximation might lead to losing some information about higher-order terms. However, it has been proven that (27) is sufficiently accurate if $\delta$ satisfies (28) [48, 37, 47], leading to the error of the approximation approaching zero. ∎
Appendix D Proof of Theorem 2
We proceed by defining $\bar{M}=M(M+1)/2$ and writing the Lagrangian of (34) as
It is obvious that by forcing $[\boldsymbol{\omega}]_{r}=0,\forall r\in\mathcal{D}_{\mathcal{G}}$, we have $\bar{\lambda}_{r}^{(l)}[\boldsymbol{\omega}]_{r}=0,\forall r\in\mathcal{D}_{\mathcal{G}}$. Meanwhile, for $r\notin\mathcal{D}_{\mathcal{G}}$, the derivative of the Lagrangian function is given by
We know that the solution must satisfy KKT conditions [42]. Hence, by setting the derivative to zero we can write the closed-form expression of the solution at the $l$th iteration as in (36). Then, using the complementary slackness for the dual variable $\nu_{r}^{(l)}$, and leveraging (36) and the fact that $\nu_{r}^{(l)}\geq 0,\forall r$, we have (37). Similar to Theorem 1, we introduce $\rho_{\omega}\leq 1$ to control the convergence of the solution, leading to (35) giving the solution at the end of the $l$th iteration. ∎

### v2

@@ IV-A Convergence Analysis
Fig. 3 illustrates the convergence performance of the IT-BDRIS, IT-WF, SCA-BDRIS, and SDR-BDRIS. It is seen that all algorithms iteratively converge toward a solution. Moreover, the number of required iterations increases with $M$ and $N$ as discussed in Section III-F. Moreover, Fig. 3.b-Fig. 3.d highlight that the proposed methods for alternating optimization of waveform and passive beamforming lead to convergence. Although the number of iterations in SDR-BDRIS is lower than in IT-BDRIS, each iteration of SDR-BDRIS requires solving multiple SDP problems with high complexity. At the same time, IT-BDRIS relies only on closed-form computations. Observe that the number of iterations for different setups is the same in SDR-BDRIS, but the time complexity of solving (27) drastically increases with $M$ since the hermitian matrix subspace scales with $M^{4}$.
@@ IV-B Algorithms Performance Comparison
Herein, we compare the performance of the proposed algorithms and consider also an optimal benchmark derived for $N=1$. Note that for a single-carrier system ($N=1$), problem (5) becomes equivalent to maximizing the RF power, i.e., the second-order term in $i_{dc}$, in terms of the solution. Thus, the framework introduced in [BD-RIS_Graph] serves as a benchmark when $N=1$ and is labeled as optimal in Fig. 4.a.
Fig. 4 compares the performance of the proposed algorithms for $N=1$ and $N=4$ as a function of $M$. Note that SDR-BDRIS achieves optimal performance, while SCA- and IT-BDRIS lead to relatively lower $i_{dc}$ values, but are still close to the optimal solution, and the gap increases with $M$. Moreover, for $N=4$, the same trend is observed, where SDR-BDRIS outperforms the other proposed approaches. However, as stated earlier, the complexity of SDR-BDRIS is considerably higher than that of SCA-BDRIS and IT-BDRIS, especially for large values of $M$; thus, a trade-off exists between performance and complexity, which is observable in the figures. Moreover, SCA-BDRIS loses performance due to constraint approximation, which may lead to divergence from the feasible space over iterations. Note that there is no optimal benchmark for the multi-carrier case. Fig. 5 compares the performance of the proposed algorithms as a function of $N$, showing that SDR-BDRIS achieves the best performance also in this case.
Figure 4: Average $i_{dc}$ at the ER as a function of $M$ using different algorithms for (a) $N=1$ (left) and (b) $N=4$ (right) with NLoS channel ($\alpha=0.1$), $P_{T}=50$ dBm, and a fully connected BD-RIS.
Figure 5: Average $i_{dc}$ at the ER as a function of NN using different algorithms with $\alpha=0.1$, $M=16$, $P_{T}=50$ dBm, and a fully connected BD-RIS.
@@ IV-C Waveform Behaviour & Channel Analysis
One of the key considerations of WPT with non-linear EH is the operating regime of the rectifier. Specifically, if the input RF power to the rectifier is relatively low, the RF-to-DC conversion efficiency drops significantly. On the other hand, if the RF input power is high such that the rectifier’s diode enters the breakdown region, the mathematical model for the small-signal regime is no longer valid. Thus, it is important to select $P_{T}$ properly such that the received RF power is in the proper regime [clerckx2018beneficial]. For this, we proceed with some discussions on the system’s behavior for different $P_{T}$ values.
@@
Remark 4.
The transmit power level of 50 dBm is chosen to ensure sufficient RF input at the rectifier under the single-antenna setup, which lacks spatial diversity gain. Although this power level is higher than typical practical values, it helps isolate the waveform design effects. In real-world systems with multi-antenna configurations, the required transmit power can be significantly reduced due to spatial beamforming and diversity gains.
Figure 6: The normalized gain of the channels and input signal at different sub-carriers for (a) $P_{T}=50$ dBm (top) and (b) $P_{T}=30$ dBm (bottom) for $\alpha=0.1$ (left), $\alpha=1$ (middle), and $\alpha=10$ (right). We assume a fully connected BD-RIS with $M=32$, $N=8$, and a pure NLoS ($\kappa=0$) random channel realization. Note that the bars are represented in a stacked fashion, meaning that bars with the same length represent the same normalized gain.
@@
Figure 7: The received signal in time domain for (a) $P_{T}=50$ dBm (top) and (b) $P_{T}=30$ dBm (bottom) for $\alpha=1$ (left) and $\alpha=10$ (right) with fully connected BD-RIS, $M=32$, and $N=8$ at a random channel realization.
@@ IV-D D-RIS versus BD-RIS
Herein, we compare the performance of D-RIS and fully connected BD-RIS. The BD-RIS results are obtained using the SDR-BDRIS approach, while for D-RIS, this approach can be straightforwardly modified as presented in [clerckxRISWPT].
Note that for far-field LoS, the channel is deterministic and is computed using the relative phase difference of elements. Thus, we assume that the variation among the sub-carriers is negligible such that for a given $\boldsymbol{\Theta}$, $h_{n}=h$. Hereby, (II-B) can be rewritten as $\alpha_{1}\left\lVert h\right\rVert^{2}+\alpha_{2}\left\lVert h\right\rVert^{4}$, where $\alpha_{1}$ and $\alpha_{2}$ are positive coefficients depending on the waveform and rectenna circuit. To maximize this, it is sufficient to maximize the cascade channel gain $\left\lVert h\right\rVert$, which can be achieved by constructing a diagonal $\boldsymbol{\Theta}$ such that $\boldsymbol{\Theta}]_{i,i}=e^{-j([\tilde{\mathbf{h}}_{R,n}]_{i}+[\tilde{\mathbf{h}}_{I,n}]_{i})},\forall i$. Such a scattering matrix can be designed using D-RIS; thus, D-RIS can achieve the optimal performance and BD-RIS does not introduce any additional gain. Moreover, for $N=1$, (5) becomes equivalent to maximizing the RF power, which allows for extending the findings of [BD-RIS_scattering_clerckx] to single-carrier WPT. Thus, for a channel with NLoS components, we expect that BD-RIS outperforms D-RIS even with a continuous wave.
@@
Remark 5.
Note that the performance gap might be different in the case of including power consumption-related constraints into the optimization problem in (5). Specifically, this introduces trade-offs between circuit complexity and performance in BD-RIS, leading to a change of performance depending on the power consumption [energy_BDRIS]. Notably, the fundamental Pareto tradeoff between performance and hardware complexity has been studied in [Pareto_Frontier_nerini, BD-RIS_Graph], and low-complexity optimal architectures exist, such as tree-connected and band/stem-connected, which can achieve the same performance as fully connected, but with lower hardware complexity [Pareto_Frontier_nerini, BD-RIS_Graph].
@@
V Conclusions and Future Works
We considered a BD-RIS-aided SISO WPT system with EH non-linearity. Moreover, we formulated a joint beamforming and waveform optimization problem aiming to maximize the harvested power at the ER. To address this problem, we decoupled the waveform optimization and beamforming problems and leveraged alternating optimization. We proposed an efficient iterative method for waveform optimization and four beamforming approaches, namely SDR-BDRIS, SDP-BDRIS, SCA-BDRIS, and IT-BDRIS. Furthermore, IT-BDRIS has the lowest computational complexity but at the cost of performance losses. Meanwhile, it was seen that SDR-BDRIS achieves the same performance as SDP-BDRIS with less computational complexity. The simulation results proved that the proposed algorithms converge to a local optimal solution, while the complexity scales with the number of RIS elements and the number of sub-carriers. Moreover, SDR-BDRIS achieves the optimal performance for a single-carrier system and outperforms IT-BDRIS in other scenarios. Interestingly, we proved that BD-RIS achieves the same performance as D-RIS under LoS conditions (in the absence of mutual coupling), while it outperforms D-RIS in NLoS cases. Furthermore, our findings demonstrated that the frequency selectivity of the channel impacts the way that BD-RIS shapes the cascade channel. Finally, we showed that the waveform power allocation to different sub-carriers depends on the transmit power and the rectifier’s operating regime.
As a prospect for future research, one can consider the following directions. The power consumption model of the BD-RIS, which depends on its architecture and the interconnections between different ports, can be analyzed to explore the trade-offs between complexity, power consumption, and performance. Moreover, the power consumption model can also include the complexity of the optimization algorithms as an additional factor, since the algorithms’ complexity increases with the number of connections between different ports of BD-RIS. While this work provides a first step toward understanding the benefits of BD-RIS in WPT with nonlinear energy harvesting, extending the framework to more general setups is a promising direction for future research. In particular, incorporating multi-antenna transmitters and receivers or supporting multiple energy receivers would enable spatial multiplexing and further enhance the power transfer efficiency.
Appendix A Proof of Proposition 1
@@
$\stackrel{{\scriptstyle\textnormal{(a)}}}{{\mathstrut{=}}}\mathrm{Tr}({\mathbf{h}_{I,n}}\mathbf{h}_{R,n}^{T}\boldsymbol{\Theta})=\mathrm{Tr}(\mathbf{H}_{n}\boldsymbol{\Theta})$
$\stackrel{{\scriptstyle\textnormal{(b)}}}{{\mathstrut{=}}}{\mathrm{Vec}(\mathbf{H}_{n})}^{T}\mathrm{Vec}(\boldsymbol{\Theta})\stackrel{{\scriptstyle\textnormal{(c)}}}{{\mathstrut{=}}}{\mathrm{Vec}(\mathbf{H}_{n})}^{T}\mathbf{P}\boldsymbol{\theta}=\mathbf{a}_{n}^{T}\boldsymbol{\theta}.$ (48)
Assume $\mathbf{D}$, $\mathbf{F}$, and $\mathbf{H}$ are arbitrary matrices. Hereby, (A) and (48) come from $\mathrm{Tr}(\mathbf{DFH})=\mathrm{Tr}(\mathbf{HDF})$ and $\mathrm{Tr}(\mathbf{D}^{T}\mathbf{F})=\mathrm{Vec}(\mathbf{D})^{T}\mathrm{Vec}(\mathbf{F})$, respectively. Moreover, by defining $\boldsymbol{\theta}\in\mathbb{C}^{\bar{M}\times 1}$ as the vector containing the lower/upper-triangle elements in $\mathbf{\Theta}$, one can design a permutation matrix $\mathbf{P}\in\{0,1\}^{M^{2}\times\bar{M}}$ such that $\mathbf{P}\boldsymbol{\omega}=\mathrm{Vec}(\mathbf{\Omega})$, which leads to (48). ∎
Appendix B Proof of Proposition 2
The Neumann approximation can be used to approximate the inverse of a matrix using another close matrix. Imagine $\mathbf{A}$ is a close matrix to an invertible $\mathbf{X}$, such that [mimoneumann]
@@ Appendix B Proof of Proposition 2
Then, the inverse of $\mathbf{A}$ can be written using the Neumann series as [stewart1998matrix]
@@
Now, let us consider $\mathbf{\Omega}$ as a small increment to the impedance matrix $\mathbf{Z}$, such that $\left\lVert[\mathbf{\Omega}]_{i,m}\right\rVert\leq\delta,\forall i,m$, leading to $\mathbf{Z}+\mathbf{\Omega}$ being close to $\mathbf{Z}$. Hereby, we can write the first-order Neumann series as in (33). Note that the first-order approximation might lead to losing some information about higher-order terms. However, it has been proven that (33) is sufficiently accurate if $\delta$ satisfies (34) [direnzomutual2, direnzoclerckxmutual, stewart1998matrix], leading to the error of the approximation approaching zero. ∎
[Uncaptioned image] Amirhossein Azarbahram (Graduate Student Member, IEEE) received the B.Sc. degree in telecommunications from KN Toosi University of Technology, Iran, in 2020, and the M.Sc. degree in communications systems from Sharif University of Technology, Iran, in 2022. He is currently a Doctoral researcher at the Centre for Wireless Communications (CWC), University of Oulu, Finland. He was a visiting researcher at the Connectivity Section (CNT) of the Department of Electronic Systems, Aalborg University (AAU), Denmark, in 2024. His research interests include wireless communications and AI-enabled integrated sensing and communications.
[Uncaptioned image] Onel L. A. López (Senior Member, IEEE) is an Associate Professor (tenure track) in sustainable wireless communications engineering at the University of Oulu, Finland. He received the B.Sc. (1st class honors, 2013), M.Sc. (2017), and D.Sc. (with distinction, 2020) degree in Electrical Engineering from the Central University of Las Villas (Cuba), the Federal University of Paraná (Brazil), and the University of Oulu (Finland), respectively. In 2013-2015, he served as a telematics specialist at the Cuban telecommunications company (ETECSA). In 2020, he was a post-doctoral researcher in a joint project between the University of Oulu and Nokia Oulu, Finland. He was on a six-month research visit to Rice University and the University of Houston, Texas, US, in 2024. He is a collaborator to the 2016 Research Award given by the Cuban Academy of Sciences, a co-recipient of the 2019 and 2023 IEEE European Conference on Networks and Communications (EuCNC) Best Student Paper Award, and the recipient of both the 2020 Best Doctoral Thesis award granted by Academic Engineers and Architects in Finland TEK and Tekniska Föreningen i Finland TFiF in 2021 and the 2022 Young Researcher Award in the field of technology in Finland. He is currently an Associate Editor of the IEEE Transactions on Communications, IEEE Wireless Communications Letters, and IEEE Communications Letters. His research interests include sustainable IoT, energy harvesting, wireless RF energy transfer, wireless connectivity, machine-type communications, and cellular-enabled sensing and positioning systems.
[Uncaptioned image] Bruno Clerckx (Fellow, IEEE) is a (Full) Professor, the Head of the Communications and Signal Processing Group, and the Head of the Wireless Communications and Signal Processing Lab, within the Electrical and Electronic Engineering Department, Imperial College London, London, U.K. He received the MSc and Ph.D. degrees in Electrical Engineering from Université Catholique de Louvain, Belgium, and the Doctor of Science (DSc) degree from Imperial College London, U.K. He spent many years in industry with Silicon Austria Labs (SAL), Austria, where he was the Chief Technology Officer (CTO) responsible for all research areas of Austria’s top research center for electronic based systems and with Samsung Electronics, South Korea, where he actively contributed to 4G (3GPP LTE/LTE-A and IEEE 802.16m). He has authored two books on “MIMO Wireless Communications” and “MIMO Wireless Networks”, over 350 peer-reviewed international research papers, and 150 standards contributions, and is the inventor of 80 issued or pending patents among which several have been adopted in the specifications of 4G standards and are used by billions of devices worldwide. His research spans the general area of wireless communications and signal processing for wireless networks. He received the prestigious Blondel Medal 2021 from France for exceptional work contributing to the progress of Science and Electrical and Electronic Industries, the 2021 Adolphe Wetrems Prize in mathematical and physical sciences and the 2024 Georges Vanderlinden Prize in Electromagnetism and Telecommunications from Royal Academy of Belgium, multiple awards from Samsung, IEEE best student paper award, and the EURASIP (European Association for Signal Processing) best paper award 2022. He is a Fellow of the IEEE and the IET.
[Uncaptioned image] Marco Di Renzo (Fellow, IEEE) received the Laurea (cum laude) and Ph.D. degrees in electrical engineering from the University of L’Aquila, Italy, in 2003 and 2007, respectively, and the Habilitation à Diriger des Recherches (Doctor of Science) degree from University Paris-Sud (currently Paris-Saclay University), France, in 2013. Currently, he is a CNRS Research Director (Professor) and the Head of the Intelligent Physical Communications group with the Laboratory of Signals and Systems (L2S) at CNRS & CentraleSupelec, Paris-Saclay University, Paris, France. Also, he is Chair Professor in Telecommunications Engineering, the Director of the Centre for Telecommunications Research, and the Head of the Telecommunications Group, Department of Engineering, King’s College London, London, United Kingdom. He was a France-Nokia Chair of Excellence in ICT at the University of Oulu (Finland), a Tan Chin Tuan Exchange Fellow in Engineering at Nanyang Technological University (Singapore), a Fulbright Fellow at The City University of New York (USA), a Nokia Foundation Visiting Professor at Aalto University (Finland), and a Royal Academy of Engineering Distinguished Visiting Fellow at Queen’s University Belfast (U.K.). He is a Fellow of the IEEE, IET, EURASIP, and AAIA; an Academician of AIIA; an Ordinary Member of the European Academy of Sciences and Arts, an Ordinary Member of the Academia Europaea, and an Ordinary Member of the Italian Academy of Technology and Engineering; an Ambassador of the European Association on Antennas and Propagation; and a Highly Cited Researcher. His recent research awards include the Michel Monpetit Prize conferred by the French Academy of Sciences, the IEEE Communications Society Heinrich Hertz Award, and the IEEE Communications Society Marconi Prize Paper Award in Wireless Communications. He served as the Editor-in-Chief of IEEE Communications Letters from 2019 to 2023. Currently, he is a Voting Member of the Fellow Evaluation Standing Committee, the Chair of the Publications Misconduct Ad Hoc Committee, and the Director of Journals of the IEEE Communications Society. Also, he is on the Editorial Board of the Proceedings of the IEEE.
[Uncaptioned image] Matti Latva-aho (Fellow, IEEE) is a distinguished expert in wireless communications. He holds M.Sc., Lic.Tech., and Dr.Tech. (Hons.) degrees in Electrical Engineering from the University of Oulu, Finland, awarded in 1992, 1996, and 1998, respectively. From 1992 to 1993, he worked as a Research Engineer at Nokia Mobile Phones in Oulu before joining the Centre for Wireless Communications (CWC) at the University of Oulu. Prof. Latva-aho served as Director of CWC from 1998 to 2006 and later as Head of the Department of Communication Engineering until August 2014. He was nominated as an Academy Professor by the Academy of Finland in 2017. He is a Professor of Wireless Communications at the University of Oulu and served as Director of the National 6G Flagship Programme. He is also a Global Fellow at The University of Tokyo. In 2025, he was appointed Vice-Rector for Research at the University of Oulu for a five-year term. With an extensive portfolio of over 600 conference and journal publications, Prof. Latva-aho has significantly advanced the field of wireless communications. His contributions were recognized in 2015 when he received the prestigious Nokia Foundation Award for his groundbreaking research in mobile communications.